OpenStudy (anonymous):

Suppose a population of 160 crickets doubles in size ever month. The function f(x)=160*2^x gives the population after x months. How many crickets will there be after two years?

OpenStudy (ybarrap):

f(2) = 160 * 2^(2) = 160 * 4 = 640 Not an exponential function but it does involved exponentiation.